The GMOS-N detector issue reported on March 2 has largely been resolved by the full thermal cycle performed between March 2 and 3. The extended bright columns on amplifiers 7 and 12 are no longer present. The narrow bad column of hot pixels on amplifier 5 persists and saturates in longer science exposures. The following figure shows the hot column in an overscan-subtracted bias image (as marked by the blue arrow).
The GMOS-N detector has shown new bias features since an uncontrolled warm-up of the detector on February 26, 2022 . These features consist of a narrow bright column on amplifier 5, and broader bright columns on amplifiers 7 and 12. As a first measure to address this new issue, a full thermal cycle of the detector is being performed starting from today. GMOS will be unavailable for the next couple of nights until the thermal cycle is complete.
GMOS-S is unavailable due to the re-appearance of the CTE problem on CCD1, plus high noise structure on CCD2. Troubleshooting started on Monday and is currently ongoing.
After a thermal cycle performed over the weekend, the detector amplifier #5 is back to its normal perfromance. CCD1 CTE is also nominal, and OIWFS performance is OK after the successful repair that had been performed during September. Therefore GMOS-S is back at full capacity.
GMOS was installed on the telescope, but it is not available yet due to the issue with CCD2 on the GMOS detector array, which features enhanced noise structure. The troubleshooting will be carried away during this week.
UPDATE (Sep 03): the repair work will extend through mid-September. The reason for this is that one of the optical components of the OI probe needs to be re-attached to the pickoff arm after it was found loose. The work will also involve the corresponding optical alignment check.
Due to a mechanical failure, The GMOS-S OIWFS will be out of service at least for the rest of August. A repair plan at the CPO Instrument Lab is being worked on. As a consequence of this situation, GMOS-S will be available only with PWFS2 as guiding option.
It was recently reported that the observed counts in spectra seemed to be lower than expected on CCD1 for some programs - also, odd emission-like features (‘spikes’) were noticed in the CCD1 portion on stellar spectra that are known to be featureless. All this applies only to data acquired during the ‘bad-CTE’ periods of CCD1, i.e. between:
As Gemini South resumed nighttime operation, GMOS-S was checked out during last weekend and started operating in queue mode. To be noted is that the CCD1 CTE problem reported on Jan10, 2020 is still present, meaning that programs using the N&S mode for spectroscopy are potentialliy affected, as well as IFU programs. These programs will remain on hold, with the exception of programs where this issue has no impact on the science (e.g. region of interest outside of CCD1)